# SMURF: Continuous Dynamics for Motion-Deblurring Radiance Fields

### v1

@@ Abstract
Neural radiance fields (NeRF) has attracted considerable attention for their exceptional ability in synthesizing novel views with high fidelity. However, the presence of motion blur, resulting from slight camera movements during extended shutter exposures, poses a significant challenge, potentially compromising the quality of the reconstructed 3D scenes. While recent studies have addressed this issue, they do not consider the continuous dynamics of camera movements during image acquisition, leading to inaccurate scene reconstruction. Additionally, these methods are plagued by slow training and rendering speed. To effectively handle these issues, we propose sequential motion understanding radiance fields (SMURF), a novel approach that employs neural ordinary differential equation (Neural-ODE) to model continuous camera motion and leverages the explicit volumetric representation method for faster training and robustness to motion-blurred input images. The core idea of the SMURF is continuous motion blurring kernel (CMBK), a unique module designed to model a continuous camera movements for processing blurry inputs. Our model, rigorously evaluated against benchmark datasets, demonstrates state-of-the-art performance both quantitatively and qualitatively. Source code will be publicly available at https://github.com/Jho-Yonsei/SMURF.
Keywords:
Neural Rendering View Synthesis Motion Deblurring
@@ 1 Introduction
Reconstructing complex 3D scenes from 2D images of different views and re-rendering novel view images represent a core problem in computer vision and graphics, with significant applications in augmented reality (AR) and virtual reality (VR). Over the past few years, photo-realistic novel view synthesis has greatly advanced with the emergence of Neural Radiance Fields (NeRF) [33]. It takes 3D spatial coordinates and 2D viewing directions (i.e. 5D vector) as inputs for mapping to the radiance and volume density, where the process is parameterized through an implicit multi-layer perceptron (MLP) based model. Exploiting classical volume rendering techniques [19], NeRF integrates the output radiance and volume density along the emitted rays, making the volume rendering process fully differentiable.
@@
Generally, NeRF variants have reconstructed 3D scenes using well-captured, noise-free images as inputs along with calibrated camera parameters. However, images captured in real-world may include various types of noise (e.g. camera motion blur, finite depth-of-field), which exert negative impacts on the network to train the complex geometry of the scene. Recently, many works [31, 25, 56, 40, 26] have been conducted on NeRF that takes camera motion-blurred images as input, which is caused by camera movement during the exposure time. Deblur-NeRF [31] first proposes a method that models blurring kernel by imitating the deconvolution method for blind image deblurring, to reconstruct 3D scenes from motion-blurred images and render sharp novel view images. Since the inception of Deblur-NeRF, various methods [25, 56, 40, 26] for estimating the blurring kernel have been proposed, but they have following limitations: a) their blurring kernel does not continuously model the camera movement during exposure time as shwon in Fig. 1. Since actual camera movement is normally continuous, it can be represented as a non-linear continuous function over time, and such continuous modeling allows for a more precise tracking of the camera movement path, even when the movement is complex or irregular. Previous works all use a method where the warped rays of the blurring kernel are computed jointly, making the kernel for images with complex camera motion imprecise. b) None of them overcome the tradeoff between performance, and the speed of training and rendering. DP-NeRF [25] shows performance improvement but suffers from slow training and rendering, whereas PDRF [40] shows fast training but offers relatively lower performance as shown in Fig. 2.
@@
Figure 1: Comparison of joint kernel generation and our sequential kernel generation. A blurry image $I_{blur}$ is acquired as the camera moves over the exposure time ($t_{0}\sim t_{\mathcal{N}}$), with images $I_{t}$ captured at each camera pose being composited together. Previous methods generate warped rays of the blurring kernel in a single step by passing through a parameterized network $h_{p}$ without considering the temporal sequence of camera motion. However, our approach iteratively estimates warped rays along the sequential camera motion trajectory.
In this paper, we propose a sequential motion understanding radiance fields (SMURF), which incorporates a novel continuous motion blur kernel (CMBK) for modeling precise camera movements from blurry images. The CMBK estimates the small change in pose regarding camera motion through the view information of each scene and the 2D pixel location of the input ray. The output values of the CMBK are not computed jointly as in previous studies but are designed to be computed sequentially according to camera motion, as shown in Fig. 1. Particularly, recognizing that the sequential camera movements share a single continuous function, we employ a neural ordinary differential equation (Neural-ODE) [7] to ensure that the sequentially computed camera movements exhibits continuity. Specifically, to reflect the physics inherent in camera motion, we apply continuous dynamics in the latent space to CMBK, and transform the latent feature into changes in ray within the physical space. Sequentially estimating the small pose change with a light shared function implies that the CMBK does not overly depend on the learning parameters. Additionally, we introduce two regularization strategies to prevent the divergence of camera motion generated through CMBK. The first is residual momentum, which ensures that the warped ray does not deviate significantly from the previous ray, thereby preventing overfitting. The second is the output suppression loss, which also serves as a regularization method to ensure that the warped ray does not diverge significantly from the initial ray.
Figure 2: Comparison of performance on synthetic scene datasets in terms of training time. The performance comparison metrics are PSNR, SSIM, and LPIPS.
Our SMURF leverages explicit volumetric representation methods [18, 6, 34, 50] as its backbone. Specifically, the tensor factorization-based approach, Tensorial Radiance Fields (TensoRF) [6], facilitates compact and efficient 3D scene reconstruction. Considering the fact that some parts may exhibit significant blur while others are almost free of blur in a single image, incomplete blurred information and complete sharp information merge into a few adjacent voxels via blurring kernel and ensure a coherent 3D scene. Therefore, we adopt the TensoRF as our backbone as it effectively mitigates the uncertainty of information caused by motion blur and enables faster training and high quality rendering compared to previous approaches. The outstanding performance of SMURF relative to its fast speed is demonstrated in Fig. 2. While there exists a deblurring approach [26] that uses voxel-based radiance fields as a backbone and shares the common goal of fast training with ours, a key difference lies in their use of Plenoxels [59] as the backbone. Furthermore, while they adopt the different explicit representation method from ours with another goal of efficient memory consumption, we diverge in our objectives, as we aim to leverage the advantages of a 3D voxel-based method when utilizing blurry images as input.
@@
To demonstrate the effectiveness of the proposed SMURF both quantitatively and qualitatively, we conduct extensive experiments on synthetic and real-world scenes. Our experimental results elucidate the advantages of the CMBK in comparison to previously presented blurring kernels.
Our main contributions are summarized as follows:
@@
    We propose a continuous motion blur kernel (CMBK) to sequentially estimate precise, continuous camera motion from blurry images.
  • $\bullet$
    We propose two regularization strategies that guide the warped rays to prevent divergence: residual momentum, and output suppression loss.
  • $\bullet$
    We significantly alleviate the tradeoff between performance and training speed exploiting CMBK and 3D tensor factorization-based representation.
@@
2 Related Work
2.1 Neural Radiance Fields (NeRF)
The synthesis of photo-realistic novel views from images of different viewpoints has attracted considerable attention with the advancement of NeRF [33]. In particular, NeRF uses 3D spatial coordinates and 2D viewing directions to map radiance and volume density, a process facilitated by implicit neural representation (INR). Following the success of NeRF, various sub-domains considering real-world scenarios have been explored. Many recent studies have extensively applied NeRF across various fields, including dynamic 3D scene modeling [37, 28, 43, 54, 38, 27, 30, 61, 3], scene relighting [48, 2, 32, 41, 53], and 3D reconstruction [55, 51, 57]. Furthermore, recent studies have applied NeRF for deblurring 3D scenes [31, 25, 40, 56], assuming blurry input images from real-world conditions and aiming to produce clean images. Additionally, due to the slow rendering caused by the costly MLP layers required to evaluate each pixel’s density and color, some studies [23, 42, 36, 6, 34, 59, 50] design networks for faster training and rendering. To address these challenges, various voxel-based explicit rendering methods [6, 34, 59, 50] have been proposed, offering competitive performance to NeRF while significantly reducing training and rendering times. We adopt an explicit representation method, TensoRF [6] as the backbone, which facilitates fast training and achieves accurate and efficient novel-view rendering.
2.2 NeRFs for Blurry Images
Recently, several studies have been conducted to deblur 3D scenes and synthesize clean novel view images using blurry input images from different views. Deblur-NeRF [31] proposes, for the first time, an implicit blurring kernel in 3D space inspired by blind deblurring methods for 2D vision [22, 17, 52, 35, 10]. This blurring kernel is intentionally trained close to the pixels of the blurry image, and during rendering, a clean image is obtained without the trained kernel. DP-NeRF [25] models a rigid kernel that assumes physical scene priors when blurry images are captured through a camera, utilizing only the scene’s view information regardless of pixel location. BAD-NeRF [56], assuming a very short camera exposure time, designs a kernel that linearly interpolates the camera motion trajectory in SE(3) space. However, these approaches estimate the change in camera pose jointly, not modeling the fact that light traces left by camera motion during exposure time appear sequentially. Therefore, we aim to design a kernel where the elements of the blurring kernel, namely the small changes in camera pose, are estimated sequentially over time.
2.3 Neural Ordinary Differential Equations (Neural ODEs)
Neural-ODEs [7] are proposed to interpret neural networks within the framework of ordinary differential equations, representing the underlying dynamics inherent in hidden parameters. Neural-ODEs model a parameterized time-continuous latent state and output a unique solution of the integral of continuous dynamics, utilizing given initial values and various numerical differential equation solvers (e.g. Euler’s method [12], Runge-Kutta method [24], Dormand-Prince-Shampine method [11]). They are extensively used to continuously connect the latent space embedded in videos [20, 39, 13, 16] or model the continuous dynamics of irregularly sampled time-series data [44]. Inspired by the fact that camera motion is continuous, we exploit Neural-ODEs to create a precise blurring kernel that models the hidden continuous dynamics inherent in camera motion.
3 Method
In this section, we detail the optimization process of SMURF. In Sec. 3.1, we explain the preliminaries about the 3D tensor factorization-based rendering method and 3D scene blind deblurring for our methodology. Sec. 3.2 explains how continuous dynamics are applied to our CMBK to generate warped rays. Sec. 3.3 discusses the objective function and the optimization process. The framework of SMURF is illustrated in Fig. 3.
3.1 Preliminary
@@
We follow the architecture of TensoRF [6], an explicit voxel-based volumetric representation utilizing CANDECOMP/PARAFAC (CP) [4, 14] and block term decomposition [9], which models an efficient view-dependent sparse voxel grid. TensoRF optimizes two 3D grid tensors, $\mathcal{G}_{\sigma},\mathcal{G}_{\mathbf{c}}\in\mathbb{R}^{FXYZ}$, for estimating volume density and view-dependent appearance feature, employing the vector-matrix (VM) decomposition that effectively extends CP decomposition. Exploiting this method, the grid (3rd-order) tensor is decomposed into low-rank 1D vectors and 2D matrices across three modes, reducing the space complexity from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n^{2})$ compared to conventional explicit representations [59]. This grid is represented as follows:
$\mathcal{G}=\sum_{r=1}^{R_{1}}\mathbf{b}_{r}^{F_{1}}\circ\mathbf{v}_{r}^{X}\circ\mathbf{M}_{r}^{YZ}+\sum_{r=1}^{R_{2}}\mathbf{b}_{r}^{F_{2}}\circ\mathbf{v}_{r}^{Y}\circ\mathbf{M}_{r}^{XZ}+\sum_{r=1}^{R_{3}}\mathbf{b}_{r}^{F_{3}}\circ\mathbf{v}_{r}^{Z}\circ\mathbf{M}_{r}^{XY},$ (1)
where $\circ$ denotes the outer product; $R_{1},R_{2}$, and $R_{3}$ indicates the number of low-rank components. $\mathbf{b}^{F}\in\mathbb{R}^{F}$ and $\mathbf{v}^{X}\in\mathbb{R}^{X}$are the vectors along the $F$ and $X$ axes in each mode, and $\mathbf{M}^{YZ}\in\mathbb{R}^{YZ}$ is the matrix in the $XY$ plane for each mode. Once the grids are defined, the radiance field at a 3D point $\mathbf{x}$ for computing volume density $\sigma$ and color $\mathbf{c}$ is defined as follows:
@@
$\mathbf{c}_{p}(\mathbf{r})=\sum_{i=1}^{N}T_{i}\left(1-\exp\left(-\sigma_{i}\delta_{i}\right)\right)\mathbf{c}_{i},~{}T_{i}=\exp\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right),$ (3)
where $T_{i}$ is the transmittance; $N$ and $\delta$ denote the number of sampled points and the step size between adjacent samples on ray $\mathbf{r}$, respectively.
Figure 3: Method overview of the SMURF. The CMBK encoder $\mathcal{E}_{\theta}$ transforms the embedded 2D pixel location $\mathbf{p}$ of the initial ray and view index into a latent feature. This feature is extended into an IVP with a parameterized derivative function $f_{\phi}$ in the latent space. Then, it is sequentially solved using a Neural-ODE along with given time $t$ and a chrono-view embedding function $\mathbf{\Psi}$, obtaining latent features for all warped rays. These features are transformed into changes of the ray (i.e., $\Delta\mathbf{p}$ and $\Delta\mathbf{o}$) through a decoder $\mathcal{D}_{\xi}$, and we get the warped rays by applying Eq. 11 to the initial ray. These rays are rendered into 2D pixel colors through a 3D grid-based method, and a blurry color is acquired by summing up the colors with weights from the decoder.
Blind Deblurring for 3D Scene.
For image blind deblurring [58, 5, 49], the blurring kernel $h$ is estimated in a fully unsupervised manner, and this process is achieved by convolving $h$ with sharp images. Ma et al. [31] apply this algorithm to NeRF, modeling an adaptive kernel $h_{p}$ for each ray,
$\mathbf{c}_{blur}(\mathbf{r})=\mathbf{c}_{p}(\mathbf{r})*h_{p}(\mathbf{r}),$ (4)
where $\mathbf{c}_{blur}$ and $\mathbf{c}_{p}$ respectively denote the color of the blurry pixel and the sharp pixel; *** indicates the convolution operator.
The conventional image deblurring kernel takes a fixed grid of size $K\times K$ centered around the location $p$. However, for 3D scene deblurring, applying such a kernel based on NeRF requires computing $\mathbf{c}_{p}$ for $K^{2}$ rays, which is inefficient in terms of memory and training time. To optimize a kernel for the 3D scene, it is necessary to produce a sparse kernel with fewer rays. Ma et al. [31] designs the sprase kernel to acquire blurry color, and for the temporally continuous camera motion, we extend the process as follows:
$\mathbf{c}_{blur}(\mathbf{r})=\sum_{i=0}^{\mathcal{N}}w_{p}^{t_{i}}\mathbf{c}_{p}(\mathbf{r}^{{t}_{i}}),~{}\text{w.r.t.}~{}\sum_{i=0}^{\mathcal{N}}w_{p}^{t_{i}}=1,$ (5)
where $\mathcal{N}$ and $t_{i}$ respectively denote the number of warped rays in camera motion and the instantaneous time during exposure; $w_{p}$ is the corresponding weight at each ray’s location, and $\mathbf{r}^{t_{i}}$ represents the warped ray due to the camera movement. By setting the number of $\mathcal{N}$ to be less than $K^{2}$, more efficient training is feasible. As the warped rays are determined by parameterized learning, the blurring kernel is deformable, not in a fixed grid manner.
@@ 3.2 Continuous Motion Blur Kernel
Camera-motion blur in images arises from camera movement during exposure time. Such movement, often due to the motion of the hand holding the camera, allows for the representation of camera pose changes as a temporally continuous function, as shown in Fig. 1. Therefore, our goal is to model the continuous movement of the camera. To assign this movement with continuity, we transform the given information of the rays into latent features and apply them to a Neural-ODE [7] to design a continuous latent space.
We embed information about the initial ray of camera motion into the latent space using a parameterized encoder $\mathcal{E}_{\theta}$, utilizing the scene’s view index $idx_{view}$ and 2D pixel location $\mathbf{p}$:
@@
where $\mathbf{z}_{t_{0}}\in\mathbb{R}^{d}$ is the latent feature of initial ray with hidden dimension $d$, and $\psi_{v}$ and $\psi_{p}$ denote embedding functions for view and pixel information, respectively. Within a small step limit of the latent feature $\mathbf{z}(t)$, the local continuous dynamic is expressed as $\mathbf{z}(t+\epsilon)=\mathbf{z}(t)+\epsilon\cdot d\mathbf{z}/dt$. To apply continuous dynamics to $\mathbf{z}(t)$, we model a parameterized neural derivative function $f$ in the latent space. The derivative of the continuous function is defined as follows:
@@
where $\phi$ denotes the learnable parameters of $f$. With $\mathbf{z}(t_{0})$ and the derivative function $f$, we define an initial value problem (IVP), and the features of subsequential rays in the latent space are obtained by the integral of $f$ from $t_{0}$ to the desired time. This dynamic leads to the format of ODE, and the process of obtaining latent features of the camera motion trajectory using various ODE solvers [12, 24, 11] is expressed as follows:
@@
Motion-blurred images all have different exposure times, but the given images lack information about them. Assuming a common exposure time for all images when obtaining latent features through the ODE solver may lead to suboptimal local minima since all the images exhibit varying degrees of blur due to different exposure times. Therefore, inspired by the difference in exposure time for images of a single scene, we define a chrono-view embedding function $\mathbf{\Psi}$ that simultaneously embeds given time $t$ and view index $idx_{view}$. Then, $f$ is expressed as follows:
@@
Exploiting $\mathbf{\Psi}$, when the conditions of $t_{n}$ and $idx_{view}$ are given to the ODE solver for the IVP, the subsequential latent feature is defined as an unique solution by Picard’s existence theorem [29].
@@
Latent features of the camera motion trajectory are predicted by CMBK and are expected to be optimized sequentially. In this process, predicted origin and direction of the ray $\mathbf{r}^{t}$ may diverge from the initial ray $\mathbf{r}$, potentially leading to a suboptimal kernel. Ma et al. [31] prevents ray divergence by applying a hyperbolic tangent function to $\Delta\mathbf{p}^{t}$ for regularization. However, such regularization is inefficient due to the varying intensity of motion blur across all the images. To address this issue, we apply a residual term to the latent derivative function $f$, implementing regularization that ensures the predicted ray $\mathbf{r}^{t_{i}}$ does not significantly deviate from the previous ray $\mathbf{r}^{t_{i-1}}$:
@@
where the $\Lambda$ is a shallow MLP for regularization and $f_{\phi}(\mathbf{z})$ is the simplified version of $f(\mathbf{z},t,\mathbf{\Psi}(t;idx_{view});\phi)$. This approach prevents the divergence of the camera trajectory, allowing rays to be warped regardless of the motion blur intensity of the image. Note that while the proposed residual momentum is implemented similarly to the methodology of residual connections in ResNet [15], the underlying conceptual frameworks of the two approaches are distinct.
@@
We encode the ray and the view information into the latent space using $\mathcal{E}_{\theta}$, and decode it into changes of the ray in camera motion trajectory using $\mathcal{D}_{\xi}$. In this process, since latent feature the initial ray $\mathbf{z}(t_{0})$ serves as the initial value for the ODE, there should be no change in the initial ray. Therefore, we apply an output suppression loss as follows:
@@
where $\lambda$ is the weight for the loss $\mathcal{L}_{supp}$. In this process, the color weight $w_{p}^{t_{0}}$ is not included for the loss. This concept is similar to the cycle consistency in CycleGAN [62]. Minimizing the changes of initial ray to zero value, we ensure that not just the initial ray but also the warped rays do not diverge, providing a regularization effect.
3.3 Optimization
Following Sec. 3.1, our goal is to minimize the difference between the blurry pixel color $\hat{\mathbf{c}}_{blur}(\mathbf{r})$ obtained by the blurring kernel and the ground truth pixel color of the motion-blurred image $\mathbf{c}_{blur}(\mathbf{r})$. Hence, the photometric loss $\mathcal{L}_{photo}$ is represented by $\mathcal{L}_{photo}=\left\|\mathbf{c}_{blur}(\mathbf{r})-\hat{\mathbf{c}}_{blur}(\mathbf{r})\right\|_{2}^{2}$. Furthermore, the parameterized 3D voxel grids $\mathcal{G}_{\sigma}$ and $\mathcal{G}_{c}$ are regularized through the total variation [1] losses $\mathcal{L}_{TV}^{\sigma}$ and $\mathcal{L}_{TV}^{c}$, respectively. Additionally, to prevent divergence of the warped rays and ensure initial ray consistency, Eq. 13 is applied, resulting in our combined objective as follows:
$\mathcal{L}=\mathcal{L}_{photo}+\lambda_{TV}^{\sigma}\mathcal{L}_{TV}^{\sigma}+\lambda_{TV}^{c}\mathcal{L}_{TV}^{c}+\lambda_{supp}\mathcal{L}_{supp}$ (14)
where the $\lambda_{TV}^{\sigma}$ and the $\lambda_{TV}^{c}$ are the weights for $\mathcal{L}_{TV}^{\sigma}$ and $\mathcal{L}_{TV}^{c}$, respectively.

### v2

@@ Abstract
Neural radiance fields (NeRF) has attracted considerable attention for their exceptional ability in synthesizing novel views with high fidelity. However, the presence of motion blur, resulting from slight camera movements during extended shutter exposures, poses a significant challenge, potentially compromising the quality of the reconstructed 3D scenes. To effectively handle this issue, we propose sequential motion understanding radiance fields (SMURF), a novel approach that models continuous camera motion and leverages the explicit volumetric representation method for robustness to motion-blurred input images. The core idea of the SMURF is continuous motion blurring kernel (CMBK), a module designed to model a continuous camera movements for processing blurry inputs. Our model is evaluated against benchmark datasets and demonstrates state-of-the-art performance both quantitatively and qualitatively.
@@ 1 Introduction
Generally, NeRF variants have reconstructed 3D scenes using well-captured, noise-free images as inputs along with calibrated camera parameters. The training of complex geometry in 3D scenes necessitates sharp images; however, in real-world scenarios, obtaining such images may not always be feasible due to various factors. For instance, to capture a sharp image, it is essential to set the camera’s aperture to a small size, thereby increasing the depth of field. However, a smaller aperture demands a significant amount of light, which consequently leads to longer exposure times. During these extended exposure, any movement of the handheld camera results in undesirable camera motion-blurred images. Recently, many works [14, 12, 22, 17, 13] have been conducted on NeRF that takes camera motion-blurred images as input. Deblur-NeRF [14] first proposes a method that models blurring kernel by imitating the deconvolution method for blind image deblurring, to reconstruct 3D scenes from motion-blurred images and render sharp novel view images. Since the inception of Deblur-NeRF, various methods [12, 22, 17, 13] for estimating the blurring kernel have been proposed. However, their blurring kernel does not continuously model the camera movement during exposure time. Since actual camera movement is normally continuous, it can be represented as a continuous function over time, and such continuous modeling allows for a more precise tracking of the camera movement path, even when the movement is complex or irregular. Previous works model the blurring kernel without considering the sequential camera movement, making the kernel for images with complex camera motion imprecise.
@@
Figure 1: Blurring process and our sequential kernel generation. A blurry image $I_{blur}$ is acquired as the camera moves over the exposure time, with images $I_{t}$ captured at each camera pose being composited together. Our approach estimates warped rays along the continuous camera motion trajectory.
In this paper, we propose a sequential motion understanding radiance fields (SMURF), which incorporates a novel continuous motion blur kernel (CMBK) for modeling precise camera movements from blurry images. The CMBK estimates the small change in pose regarding continuous camera motion. The output values of the CMBK are not computed jointly as in previous studies but are designed to be computed sequentially according to camera motion, as shown in Fig. 1. Additionally, we adopt explicit volumetric representation methods [9, 2, 16, 21] as its backbone. Specifically, the tensor factorization-based approach, Tensorial Radiance Fields (TensoRF) [2], facilitates compact and efficient 3D scene reconstruction.
@@
To demonstrate the effectiveness of the proposed SMURF, we conduct extensive experiments on synthetic and real-world scenes. Our experimental results elucidate the advantages of the CMBK in comparison to previously presented blurring kernels. Our main contributions are summarized as follows:
@@
    We propose a continuous motion blur kernel (CMBK) to sequentially estimate continuous camera motion from blurry images mimicking the real-world motion blur.
@@
2 Preliminary
Blind Deblurring for 3D Scene.
Deblur-NeRF [14] apply the image blind deblurring algorithm to NeRF, modeling an adaptive blurring kernel for each ray to get the blurry pixel. Deblur-NeRF designs the sprase kernel to acquire blurry color, and for the temporally continuous camera motion, we extend the process as follows:
$\mathbf{c}_{blur}(\mathbf{r})=\sum_{i=0}^{\mathcal{N}}w_{p}^{t_{i}}\mathbf{c}_{p}(\mathbf{r}^{{t}_{i}}),~{}w.r.t.~{}\sum_{i=0}^{\mathcal{N}}w_{p}^{t_{i}}=1,$ (1)
where $\mathcal{N}$ and $t_{i}$ respectively denote the number of warped rays in camera motion and the instantaneous time during exposure; $w_{p}$ is the corresponding weight at each ray’s location, and $\mathbf{r}^{t_{i}}$ represents the warped ray.
@@ 2 Preliminary
We follow the architecture of TensoRF [2], an explicit voxel-based volumetric representation utilizing CP [1, 8] and block term decomposition [5], which models an efficient view-dependent sparse voxel grid. TensoRF optimizes two 3D grid tensors, $\mathcal{G}_{\sigma},\mathcal{G}_{\mathbf{c}}\in\mathbb{R}^{FXYZ}$, for estimating volume density and view-dependent appearance feature, employing the vector-matrix (VM) decomposition that effectively extends CP decomposition. The voxel grid is represented as follows:
$\mathcal{G}=\sum_{r=1}^{R_{1}}\mathbf{v}_{r}^{X}\circ\mathbf{M}_{r}^{YZ}+\sum_{r=1}^{R_{2}}\mathbf{v}_{r}^{Y}\circ\mathbf{M}_{r}^{XZ}+\sum_{r=1}^{R_{3}}\mathbf{v}_{r}^{Z}\circ\mathbf{M}_{r}^{XY},$ (2)
where $\circ$ denotes the outer product; $R_{1},R_{2}$, and $R_{3}$ indicates the number of low-rank components. $\mathbf{v}^{X}\in\mathbb{R}^{X}$ is the vector along the $X$ axes in each mode, and $\mathbf{M}^{YZ}\in\mathbb{R}^{YZ}$ is the matrix in the $YZ$ plane for each mode. Once the grids are defined, the radiance field at a 3D point $\mathbf{x}$ for computing volume density $\sigma$ and color $\mathbf{c}$ is defined as follows:
@@
$\mathbf{c}_{p}(\mathbf{r})=\sum_{i=1}^{N}T_{i}\left(1-\exp\left(-\sigma_{i}\delta_{i}\right)\right)\mathbf{c}_{i},$ (4)
where $T_{i}=\exp\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right)$ is the transmittance; $N$ and $\delta$ denote the number of sampled points and the step size between adjacent samples on ray $\mathbf{r}$, respectively.
Figure 2: Method overview of the SMURF. The latent feature of initial ray is extended into an IVP with a parameterized derivative function $f_{\phi}$. Then, it is solved by Neural-ODEs along with given time $t$, obtaining latent features for all warped rays.
3 Method
@@ 3.1 Continuous Motion Blur Kernel
Our goal is to model the continuous movement of the camera within the exposure time $(t_{0}\leq t<t_{\mathcal{N}})$. To reflect the physics inherent in camera motion, we apply continuous dynamics to the latent space of camera motion, and transform the latent feature into changes in ray within the physical space. For this process, we transform the given information of the rays into latent features and apply them to a Neural-ODEs [3] to design a continuous latent space.
As shown in Fig. 2, we embed information about the initial ray into the latent space using a parameterized encoder $\mathcal{E}_{\theta}$, utilizing the scene’s view index $idx_{view}$ and 2D pixel location $\mathbf{p}$:
@@
where $\mathbf{z}(t_{0})\in\mathbb{R}^{d}$ is the latent feature of initial ray with hidden dimension $d$, and $\psi_{v}$ and $\psi_{p}$ denote embedding functions for view and pixel information, respectively. Within a small step limit of the latent feature $\mathbf{z}(t)$, the local continuous dynamic is expressed as $\mathbf{z}(t+\epsilon)=\mathbf{z}(t)+\epsilon\cdot d\mathbf{z}/dt$. To apply continuous dynamics to $\mathbf{z}(t)$, we model a parameterized neural derivative function $f$ in the latent space. The derivative of the continuous function is defined as follows:
@@
where $\phi$ denotes the learnable parameters of $f$. With $\mathbf{z}(t_{0})$ and the derivative function $f$, we define an initial value problem (IVP), and the features of subsequential rays in the latent space are obtained by the integral of $f$ from $t_{0}$ to the desired time. This dynamic leads to the format of ODEs, and the process of obtaining latent features of the camera motion trajectory using various ODE solvers [7, 11, 6] is expressed as follows:
@@
Inspired by the difference in exposure time for images, we define a chrono-view embedding function $\mathbf{\Psi}$ with a single-layer MLP that simultaneously embeds given time $t$ and view index $idx_{view}$. Then, $f$ is expressed as follows:
@@
Latent features of the camera motion trajectory are predicted by CMBK and are expected to be optimized continuously. In this process, predicted origin and direction of the ray $\mathbf{r}^{t}$ may diverge from the initial ray $\mathbf{r}$, potentially leading to a suboptimal kernel. To address this issue, we apply a residual term to $f$, implementing regularization that ensures the predicted ray $\mathbf{r}^{t_{i}}$ does not significantly deviate from the previous ray $\mathbf{r}^{t_{i-1}}$:
@@
Figure 3: Qualitative comparisons on synthetic scenes and real-world scenes.
@@
We encode the ray information into the latent space, and decode it into changes of the ray in motion trajectory using $\mathcal{D}_{\xi}$. In this process, since initial latent feature $\mathbf{z}(t_{0})$ serves as the initial value for the ODE, there should be no change in the initial ray. Therefore, we apply an output suppression loss as follows:
@@
where $\lambda$ is the weight for the loss $\mathcal{L}_{supp}$. In this process, the color weight $w_{p}^{t_{0}}$ is not included for the loss. Minimizing the changes of initial ray to zero value, we ensure that not just the initial ray but also the warped rays do not diverge, providing a regularization effect.